Main INDEX
Monthly INDEX
PREV
NEXT
User name A. Camsonne
Log entry time 05:56:18 on August 19, 2006
Entry number 180851
keyword=Q1 back up
put 0 in the set point current to clear the errors.
Power supply would not get back on though, we hit thw green reset button
for Q1 in the middle room.
Power supply came back on after that and magnet was ramped back to 395.95
A for set current